OpenStudy (samigupta8):

Let f(x)= integral of (t^2+2t+2)dt within limits 0 to x where x is set of real numbers satisfying the inequation log(1+√6x-x^2-8)>=0. If f(x)€[a,b] then maximum value of |a-b| is

OpenStudy (samigupta8):

@freckles

OpenStudy (samigupta8):

@ikram002p

OpenStudy (anonymous):

You could start by solving the integral. Also consider: \[\log(X) \ge 0 \]is true when \[X \ge 1\] for real numbers.

OpenStudy (samigupta8):

That gives me x€[2,4]

OpenStudy (samigupta8):

For the condition of log to be greater than

OpenStudy (samigupta8):

0

OpenStudy (anonymous):

Is it sqrt(6)x or sqrt(6x) in your inequality?

OpenStudy (samigupta8):

The surd is over the whole term 6x-x^2-8 whole of if

OpenStudy (samigupta8):

It*

OpenStudy (anonymous):

x€[2,4] is correct for the inequality, yes..

OpenStudy (freckles):

since g(t)=t^2+2t+2 is an increasing function and all above the x-axis then the area in increasing as we move to the right of x=0 Since x can only be between x=2 and x=4 then that would mean that \[\int\limits_0^2 g(t) dt \le f(x) \le \int\limits_0^4 g(t) dt\] so you will find a and b by evaluating both of those integrals I believe

OpenStudy (freckles):

that is increasing function on the interval [0,4] which is all we need to be concerned with

OpenStudy (freckles):

of course it is increasing elsewhere but that is the only interval as I said we should be worried about

OpenStudy (samigupta8):

Bt when we r gonna solve the integral won't there come a constant ... How to evaluate its value?

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

@hartnn

OpenStudy (freckles):

I'm not sure what you are asking

OpenStudy (freckles):

we already know that \[f(x) \in [\int\limits_0^2 g(t) dt, \int\limits_0^4 g(t) dt]\] so you already know your a and b you just have to evaluate those integrals

OpenStudy (freckles):

g(t) was equal to your integrand g(t)=t^2+2t+2

OpenStudy (samigupta8):

It's integration is t^3/3 + t^2 + 2t +c what about the value of c?

OpenStudy (freckles):

you have a definite integral

OpenStudy (freckles):

not an indefinite

OpenStudy (samigupta8):

Oops !! Forgot that sorry for asking a stupid question as that!! Thanks also

OpenStudy (freckles):

however you could put plus C in but the C's are just going to cancel out when you do the subtraction part

OpenStudy (freckles):

\[a=[\frac{t^3}{3}+t^2+2t]_0^2 =? \\ b=[\frac{t^3}{3}+t^2+2t]_0^4=?\]

OpenStudy (samigupta8):

a=32/3 Ans b=136/3

OpenStudy (samigupta8):

And*

OpenStudy (samigupta8):

@freckles thankyou

OpenStudy (freckles):

np